OpenStudy (anonymous):

write an equation for (0,5) and (1,0)

OpenStudy (jdoe0001):

w0t?

OpenStudy (jdoe0001):

I gather you meant a line who crosses through the points of (0,5) and (1,0)

OpenStudy (anonymous):

they are asking for the equation the y=mx+b

OpenStudy (jdoe0001):

get the slope of the line from the points given then once you have the slope, use either point and plug it along with the slope found, in the \(\bf \text{point-slope form}\) of the equation \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(0\quad ,&5)\quad &(1\quad ,&0) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"} \)